OpenStudy (punderson):

@Mehek14 The table below shows the values of f(n) for different values of n: n 1 2 3 4 5 6 f(n) 1 2 5 12 29 70 Which recursive function best represents the values shown in the table? (4 points) f(1) = 1, f(2) = 2, f(n) = 2f(n -1) f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = 2f(n -1) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) f(n - 2); n > 2

jimthompson5910 (jim_thompson5910):

Let's focus on choice A for now f(1) = 1, f(2) = 2, f(n) = 2f(n -1) f(n - 2); n > 2 specifically, I want to focus on the last part where it says `f(n) = 2f(n -1) f(n - 2); n > 2` Let's plug in n = 3 and see what happens \[\Large f(n) = 2f(n -1) f(n - 2)\] \[\Large f({\color{red}{n}}) = 2f({\color{red}{n}} -1) f({\color{red}{n}} - 2)\] \[\Large f({\color{red}{3}}) = 2f({\color{red}{3}} -1) f({\color{red}{3}} - 2)\] \[\Large f(3) = 2*f(2)*f(1)\] \[\Large f(3) = 2*2*1\] \[\Large f(3) = 4\] Hopefully you see how this all works?